# Generalizing quantum dimensions: Symmetry-based classification of local pseudo-Hermitian systems and the corresponding domain walls

### v1

@@ Abstract
One of the most fundamental ideas in the studies on non-Hermitian systems is the similarity transformation, which maps a pseudo-Hermitian system to the Hermitian system [14, 15, 16]. Especially, when the system has the $PT$ symmetry, this has played an important role[17]. In quantum field theory, a kind of law of thermodynamics, called Zamolodchikov’s $c$-theorem[18] has been generalized to $c_{\text{eff}}$-theorem[19] for $PT$ symmetric models where $c$ is the central charge of a CFT and $c_{\text{eff}}$ is the effective central charge. The phenomenology of this theorem is simple: the effective central charge $c_{\text{eff}}$ of a system should not increase during renormalization group (RG) flow under some suitable assumptions, such as $PT$ symmetry. Hence, one may expect a dual relation between the $c_{\text{eff}}$-theorem for non-Hermitian systems and the $c$-theorem for the corresponding Hermitian systems, but the quantum field theoretic (QFT) formulation of the similarity transformation has not been studied widely. Moreover, it is known that the similarity transformation can change the locality of the system, and this is analogous to the gauge transformations[20]. Hence, at this stage, a local pseudo-Hermitian system is more treatable than the corresponding nonlocal Hermitian system in QFT  [^1] We will study the symmetry topological field theories of nonlocal Hermitian systems constructed from local pseudo-Hermitian systems by the similarity transformations elsewhere.. In this sense, characterizations of pseudo-Hermitian systems and their quantum phase have still been under development in the QFT language. We note related earlier studies[22, 23] and recent one[24]. It is also worth noting that the corresponding level-rank duality[25, 26, 27] has been studied in [28, 29, 30] with a close connection to the 3d3d3d-3d correspondence[31, 32].
@@ III Verlinde line operator and algebraic generalized quantum dimension
where $\mathbb{Z}$ is the set of integer numbers. Corresponding to the eigenvalue of $L_{0}$, and the primary field $\alpha$ has conformal dimension $h_{\alpha}$. By applying the $\{L\}_{m\geq-1}$ recursively to the primary states, one can obtain the corresponding descendant states denoted as $|\alpha,M\rangle$.
@@
where $N$ is the nonnegative integer matrix and the symbol $\times$ represents the matrix multiplication. Hence, one can say that the Verlinde line is a typical object satisfying the algebraic relation between anyons in a CFT, called the fusion ring. This algebraic aspect has captured attentions in the field in the term, symmetry topological field theory (SymTFT)[117, 118, 119], by combining these algebraic data and Moore-Seiberg data[75, 76, 77]. We also note the related earlier work [120] which contains the basic idea of the sandwich construction in the recent terminolgy[121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 66, 131]. We stress that the fusion ring $\{\mathcal{Q}_{\alpha}\}$ is defined as a ring over $\mathbb{C}$. More phenomenologically, $\{\mathcal{Q}_{\alpha}\}$ should be treated as a set of conserved charges that has a close connection to topological defects forming a fusion category. Historically, this interpretation of conserved charges has appeared in [39], to demonstrate the physical realization of fusion category structure studied in earlier mathematical literature. In other words, in those days, conserved charges realizing rings over $\mathbb{C}$ were thought of as more physical than category theory. We also note that the distinction between conserved charges and topological defects plays a role in studying corresponding lattice realizations[132, 133, 134, 135, 136, 137].
We also note that the noninvertible symmetry operator defined by linear algebra (or zero of the modular $S$ matrix in our case) can be different from the “noninvertible symmetry" defined in category theories [^7] For example, the Fibonacci anyon $\tau$ with the fusion rule $\tau\times\tau=I+\tau$ generates invertible symmetry $\mathcal{Q}_{\tau}$ under the relation $\mathcal{Q}_{\tau}(\mathcal{Q}_{\tau}-\mathcal{Q}_{I})=\mathcal{Q}_{I}$.. In other words, there can exist categorically noninvertible symmetry operators that are invertible in linear algebra. However, when studying quantum Hamiltonian formalism based on linear algebra, the definition in linear algebra is more dominant when studying the RGs. We believe that there exists some categorical formalisms consistent with the observations in linear algebra, but they seem less common in the physics literature. To avoid confusion, we notify the “noninvertible symmetry" in category theory as categorical noninvertible symmetry or non-group-like symmetry. It also seems reasonable to call the corresponding objects nonabelian symmetry when considering its connection to nonabelian anyons in physics, but the term “nonabelian” also appears in group theory and the corresponding group rings in a different meaning.
@@
where $\alpha$ represent type of symmetry or anyons, $a$ represents sector. (We note that our generalization is different from the generalized quantum dimension introduced in the study of representation theory of Grassmannian[139].) By choosing $a$ as the label of the lowest energy state $o$ or the effective vacuum, it is known that the above become positive numbers in unitary CFTs and a wide class of nonunitary CFTs[140, 141]. In a unitary CFT, the relation $o=I$ holds, but in a nonunitary CFT, this does not hold. This swapping of the role of the exact vacuum $I$ and the effective vacuum $o$ is called Galois shuffle[140] and played a significant role in proving modular noninvariance of gapless fractional quantum Hall states[142, 143, 144, 98, 66]. We also note that the quantity $q_{\alpha,(a)}$ has already appeared in a few mathematical studies [145, 146], or as the coefficients of the Verlinde operators[39], as we revisited Eq. (7). However, its implications for RG flows have not been studied to our knowledge. One can also introduce the antichiral analog of the algebraic generalized quantum dimensions $\overline{q_{\alpha,(a)}}$, and the distinction coming from the chirality plays a fundamental role in studying more general models, such as $Z_{N}$ symmetric models[147, 66, 48, 131]. However, for simplicity, we mainly focus on the $Z_{2}$ symmetric cases, and focus on the chiral structures in theories. Problems resulting from the chiralities of theories are discussed in Sec. VI.
@@
The boundary entropy is defined by evaluating the dominant contributions of the limit $it\rightarrow$-$\infty$. From the open string basis, this is a high temperature limit and difficult to evaluate directly, but from the closed string basis, it can be evaluated easily by observing the coefficient before the lowest energy states. Hence, one can obtain the following dominant asymptotics,
@@ IV Massless flows and matching algebraic generalized quantum dimensions
This relation also implies that the objects in $|s_{1}|$ are noninvertible by assuming the matrix representation of $\{\mathcal{Q}\}$ and their linear algebra. In a similar way, one can construct multiple $(a)$ sectors preserving homomorphism, and we denote the preserved sector as $\mathbb{S}_{\rho}$. Hence the following defines the $\mathbb{S}_{\rho}$ preserving ideal,
@@
where we have used ’ to indicate the objects in the IR theories, and $\mathbb{S}^{\prime}_{\rho,\times}$ is the pair of the preserved sector from UV to IR. We also introduce the form of the ring homomorphism as follows,
@@ V.2 $M(3,4)$ to $M(2,5)$
$\rho(\psi)$ $=A_{I^{\prime}}^{\psi}I^{\prime}+A_{\tau^{\prime}}^{\psi}\tau^{\prime},$ (72)
@@ V.4 $M(2,7)$ to $M(2,5)$
$\begin{split}&\quad 125(A_{\Phi}^{I^{\prime}})^{6}-250(A_{\Phi}^{I^{\prime}})^{5}-25(A_{\Phi}^{I^{\prime}})^{4}\\ &+225(A_{\Phi}^{I^{\prime}})^{3}-30(A_{\Phi}^{I^{\prime}})^{2}-45A_{\Phi}^{I^{\prime}}+1=0.\end{split}$ (105)
@@
In conclusion, we find six solutions satisfying the homomorphism conditions. Interestingly, there exists only one positive solution $(A_{\Phi}^{I^{\prime}},A_{\Phi}^{\tau^{\prime}},A_{\Psi}^{I^{\prime}},A_{\Psi}^{\tau^{\prime}})=(0.820079,0.606822,0.0407623,1.36352)$. We conjecture that this positive solution corresponds to the existing massless RG flow, and further studies based on other methods are necessary for the confirmation of our prediction. From the analysis so far, we observe that these conditions lead to algebraic equations such as Eq. (105). It should be noted that the identification of the massless RGs as projections has been proposed in the study of $M(2,2k+1)$ models [238].
@@ VI Spin-classification for non-group-like objects: (Half) integer spin conditions
Interestingly, there exists a nontrivial structure when restricting our attention to the chiral and antichiral property of the flow $M(4,5)\rightarrow M(3,4)$ preserving the Ising fusion ring, $\{I,\psi,\sigma\}\sim\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$. Assuming the $Z_{2}$ symmetry preservation, one can consider RG flows $\{I,\psi\}\rightarrow\{I,\psi^{\prime}\}$ and $\{I,\psi\}\rightarrow\{\overline{I}^{\prime},\overline{\psi}^{\prime}\}$ because these the $Z_{2}$ simple currents satisfying the $Z_{2}$ anomaly condition $h_{\psi}-h_{\psi^{\prime}}=0\ (\text{mod}.1)$ or $h_{\psi}+\overline{h}_{\overline{\psi}^{\prime}}=0\ (\text{mod}.1)$. This anomaly matching condition, a kind of ’t Hooft anomaly matching codition[239], is nothing but Lieb-Schultz-Mattis anomaly matching condition[240, 241, 242] with a close connection to the Haldane conjecture[243, 244, 245, 246, 247], and this condition results in the appearance of the general electron operators constructing the singlevalued wavefunction of the TQFT $\mathbf{A}\otimes\mathbf{A}^{\prime}$ or $\mathbf{A}\otimes\overline{\mathbf{A}}^{\prime}$ [204](or its $Z_{2}$ extension more precisely[248, 249, 250, 251, 252]). Hence, one will expect the same condition for the preserved structure $\{I,\psi,\sigma\}\sim\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$, but $\sigma$ and $\sigma^{\prime}$ only satisfies the relation $h_{\sigma}+h_{\sigma^{\prime}}=0$. Hence, the flow $\{I,\psi,\sigma\}\rightarrow\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$ only preserves $Z_{2}$ structure when applying the anomaly classification to the non-group like symmetry whereas the flow $\{I,\psi,\sigma\}\rightarrow\{\overline{I}^{\prime},\overline{\psi}^{\prime},\overline{\sigma}^{\prime}\}$ is an anomaly preserving flow. More categorically, these two flows correspond to the choice of the common connected etale algebra $\{I,\psi\}=\{I^{\prime},\psi^{\prime}\}$ or $\{I,\psi,\sigma\}=\{\overline{I}^{\prime},\overline{\psi}^{\prime},\overline{\sigma}^{\prime}\}$. Consequently, if one believes that the massless flow $M(4,5)\rightarrow M(3,4)$ preserves the Ising fusion ring structure in the anomaly-free way, the application of the corresponding domain wall should change the chirality of objects.
@@ VIII Acknowledgement
The authors thank Po-Yao Chang, Naomichi Hatano, and Hosho Katsura for sharing their knowledge on non-Hermitian systems. YF thanks Yuma Furuta and Shinichiro Yahagi for related collaborations. YF also thanks Sylvain Ribault and Ingo Runkel for the discussions on fusion rules and Jurgen Fuch and Yuji Tachikawa for the helpful discussions on the distinction between fusion categories and fusion rings. We thank Ingo Runkel and Gerard Watts for helpful comments and for checking some technical calculations in the manuscript. Also, we thank Tadashi Takayanagi for commenting on topological entanglement entropy and its importance. YF thanks the support from NCTS. TK is supported by Grant-in-Aid for JSPS Fellows No. 23KJ1315.
@@ A.2 Phenomenology on entanglement cut: Naturality of complex spectrum
First thing to note is that the entanglement spectrum in a pseudo-Hermitian system can be complex[269, 270, 271, 272]. However, following the discussion in the previous sections, this phenomenon seems unusual because the complex energy value cannot appear in the previous discussion. Moreover, there exist established studies on the universal behaviors of real entanglement entropy in the nonunitary CFTs by analyzing the corresponding two dimensional statistical models and their quantum group structure[273, 274], combined with the twist field method[275, 276, 277]. We also note [278] as a reference summarizing a variant of related entropies, such as those in [279, 280].

### v2

@@ Abstract
One of the most fundamental ideas in the studies on non-Hermitian systems is the similarity transformation, which maps a pseudo-Hermitian system to the Hermitian system [14, 15, 16]. Especially, when the system has the parity-time (PT) symmetry, this has played an important role[17]. In quantum field theory, a kind of law of thermodynamics, called Zamolodchikov’s $c$-theorem[18] has been generalized to $c_{\text{eff}}$-theorem[19] for $PT$ symmetric models where $c$ is the central charge of a CFT and $c_{\text{eff}}$ is the effective central charge. The phenomenology of this theorem is simple: the effective central charge $c_{\text{eff}}$ of a system should not increase during renormalization group (RG) flow under some suitable assumptions, such as $PT$ symmetry. Hence, one may expect a dual relation between the $c_{\text{eff}}$-theorem for non-Hermitian systems and the $c$-theorem for the corresponding Hermitian systems, but the quantum field theoretic (QFT) formulation of the similarity transformation has not been studied widely. Moreover, it is known that the similarity transformation can change the locality of the system, and this is analogous to the gauge transformations[20]. Hence, at this stage, a local pseudo-Hermitian system is more treatable than the corresponding nonlocal Hermitian system in QFT (For readers interested in the implications of the similarity transformation, we note the complementary work by the authors [21] and the references therein). In this sense, characterizations of pseudo-Hermitian systems and their quantum phase have still been under development in the QFT language. We note related earlier studies[22, 23] and recent one[24]. It is also worth noting that the corresponding level-rank duality[25, 26, 27] has been studied in [28, 29, 30] with a close connection to the 3d3d3d-3d correspondence[31, 32].
@@ III Verlinde line operator and algebraic generalized quantum dimension
where $\mathbb{Z}$ is the set of integer numbers. Corresponding to the eigenvalue of $L_{0}$, and the primary field $\alpha$ has conformal dimension $h_{\alpha}$. By applying the $\{L\}_{m\leq-1}$ recursively to the primary states, one can obtain the corresponding descendant states denoted as $|\alpha,M\rangle$.
@@
where $N$ is the nonnegative integer matrix and the symbol $\times$ represents the matrix multiplication. Hence, one can say that the Verlinde line is a typical object satisfying the algebraic relation between anyons in a CFT, called the fusion ring. This algebraic aspect has captured attentions in the field in the term, symmetry topological field theory (SymTFT)[117, 118, 119], by combining these algebraic data and Moore-Seiberg data[75, 76, 77]. We also note the related earlier work [120] which contains the basic idea of the sandwich construction in the recent terminolgy[121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 66, 131]. We also note that the application of SymTFTs to a class of nonequilibrium phenomena, Floquet systems, can be seen in [132, 133]. We stress that the fusion ring $\{\mathcal{Q}_{\alpha}\}$ is defined as a ring over $\mathbb{C}$. More phenomenologically, $\{\mathcal{Q}_{\alpha}\}$ should be treated as a set of conserved charges that has a close connection to topological defects forming a fusion category. Historically, this interpretation of conserved charges has appeared in [39], to demonstrate the physical realization of fusion category structure studied in earlier mathematical literature. In other words, in those days, conserved charges realizing rings over $\mathbb{C}$ were thought of as more physical than category theory. Related to this aspect, it is known that the distinction between conserved charges and topological defects plays a role in studying corresponding lattice realizations[134, 135, 136, 137, 138, 139].
We also note that the noninvertible symmetry operator defined by linear algebra (or zero of the modular $S$ matrix in our case) can be different from the “noninvertible symmetry" defined in category theories [^6] For example, the Fibonacci anyon $\tau$ with the fusion rule $\tau\times\tau=I+\tau$ generates invertible symmetry $\mathcal{Q}_{\tau}$ under the relation $\mathcal{Q}_{\tau}(\mathcal{Q}_{\tau}-\mathcal{Q}_{I})=\mathcal{Q}_{I}$.. In other words, there can exist categorically noninvertible symmetry operators that are invertible in linear algebra. However, when studying quantum Hamiltonian formalism based on linear algebra, the definition in linear algebra is more dominant when studying the RGs. We believe that there exists some categorical formalisms consistent with the observations in linear algebra, but they seem less common in the physics literature [^7] A $\mathbb{C}$-linear category will correspond to the linear algebra, but it is less common in the context of fusion rules. To avoid confusion, we notify the “noninvertible symmetry" in category theory as categorical noninvertible symmetry or non-group-like symmetry. It also seems reasonable to call the corresponding objects nonabelian symmetry when considering its connection to nonabelian anyons in physics, but the term “nonabelian” also appears in group theory and the corresponding group rings in a different meaning.
@@
where $\alpha$ represents type of symmetry or anyons, $a$ represents sector. (We note that our generalization is different from the generalized quantum dimension introduced in the study of representation theory of Grassmannian[142].) By choosing $a$ as the label of the lowest energy state $o$ or the effective vacuum, it is known that the above become positive numbers in unitary CFTs and a wide class of nonunitary CFTs[143, 144]. In a unitary CFT, the relation $o=I$ holds, but in a nonunitary CFT, this does not hold. This swapping of the role of the exact vacuum $I$ and the effective vacuum $o$ is called Galois shuffle[143] and played a significant role in proving modular noninvariance of gapless fractional quantum Hall states[145, 146, 147, 98, 66]. We also note that the quantity $q_{\alpha,(a)}$ has already appeared in a few mathematical studies [148, 149], or as the coefficients of the Verlinde operators[39], as we revisited Eq. (7). However, its implications for RG flows have not been studied to our knowledge. One can also introduce the antichiral analog of the algebraic generalized quantum dimensions $\overline{q_{\alpha,(a)}}$, and the distinction coming from the chirality plays a fundamental role in studying more general models, such as $Z_{N}$ symmetric models[150, 66, 48, 131]. However, for simplicity, we mainly focus on the $Z_{2}$ symmetric cases, and focus on the chiral structures in theories. Problems resulting from the chiralities of theories are discussed in Sec. VI with a possible connection to the Alice ring in literature[151, 152].
@@
The boundary entropy is defined by evaluating the dominant contributions of the limit $it\rightarrow$-$\infty$. From the open string basis, this is a high temperature limit and difficult to evaluate directly, but, from the closed string basis, it can be evaluated easily by observing the coefficient before the lowest energy states. Hence, one can obtain the following dominant asymptotics,
@@ IV Massless flows and matching algebraic generalized quantum dimensions
This relation also implies that the objects in $|s_{1}|$ are noninvertible by assuming the matrix representation of $\{\mathcal{Q}\}$ and their linear algebra. In a similar way, one can construct homomorphisms preserving multiple $(a)$ sectors, and we denote the preserved sector as $\mathbb{S}_{\rho}$. Hence the following defines the $\mathbb{S}_{\rho}$ preserving ideal,
@@
where we have used ’ to indicate the objects in the IR theories, and $\mathbb{S}_{\rho,\times}$ is the pair of the preserved sector from UV to IR. We also introduce the form of the ring homomorphism as follows,
@@ V.2 $M(3,4)$ to $M(2,5)$
$\rho(\psi)$ $=A^{I^{\prime}}_{\psi}I^{\prime}+A^{\tau^{\prime}}_{\psi}\tau^{\prime},$ (72)
@@ V.4 $M(2,7)$ to $M(2,5)$
$\begin{split}&\quad 125(A_{\Phi}^{I^{\prime}})^{6}-250(A_{\Phi}^{I^{\prime}})^{5}-25(A_{\Phi}^{I^{\prime}})^{4}\\ &+225(A_{\Phi}^{I^{\prime}})^{3}-30(A_{\Phi}^{I^{\prime}})^{2}-45A_{\Phi}^{I^{\prime}}-1=0.\end{split}$ (105)
@@
In conclusion, we find six solutions satisfying the homomorphism conditions. Interestingly, there exists only one positive solution $(A_{\Phi}^{I^{\prime}},A_{\Phi}^{\tau^{\prime}},A_{\Psi}^{I^{\prime}},A_{\Psi}^{\tau^{\prime}})=(0.820079,0.606822,0.0407623,1.36352)$, and the resultant homomorphism has the preserving sectors $\mathbb{S}_{\rho,\times}=\{(I,I^{\prime}),(o,o^{\prime})\}$. We conjecture that this positive solution corresponds to the existing massless RG flow, and further studies based on other methods are necessary for the confirmation of our prediction. Moreover, there exists a controversy between the perturbative expressions on the flow $M(2,7)\rightarrow M(2,5)$ in [243, 244] and those in [245], and this controversy will be explained by the nonuniqueness of the homomorphism in this subsection [^12] We thank Andrei Katsevich for notifying us the subtelties of $M(2,7)\rightarrow M(2,5)$ in the literature.. From the analysis so far, we observe that these conditions lead to algebraic equations such as Eq. (105). It should be noted that the identification of the massless RGs as projections has been proposed in the study of $M(2,2k+1)$ models [247].
@@ VI Spin-classification for non-group-like objects: (Half) integer spin conditions
Interestingly, there exists a nontrivial structure when restricting our attention to the chiral and antichiral property of the flow $M(4,5)\rightarrow M(3,4)$ preserving the Ising fusion ring, $\{I,\psi,\sigma\}\sim\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$. Assuming the $Z_{2}$ symmetry preservation, one can consider RG flows $\{I,\psi\}\rightarrow\{I,\psi^{\prime}\}$ and $\{I,\psi\}\rightarrow\{\overline{I}^{\prime},\overline{\psi}^{\prime}\}$ because these the $Z_{2}$ simple currents satisfying the $Z_{2}$ anomaly condition $h_{\psi}-h_{\psi^{\prime}}=0\ (\text{mod}.1)$ or $h_{\psi}+\overline{h}_{\overline{\psi}^{\prime}}=0\ (\text{mod}.1)$. This anomaly matching condition, a kind of ’t Hooft anomaly matching codition[248], is nothing but Lieb-Schultz-Mattis anomaly matching condition[249, 250, 251] with a close connection to the Haldane conjecture[252, 253, 254, 255, 256], and this condition results in the appearance of the general electron operators constructing the singlevalued wavefunction of the TQFT $\mathbf{A}\otimes\mathbf{A}^{\prime}$ or $\mathbf{A}\otimes\overline{\mathbf{A}}^{\prime}$ [209](or its $Z_{2}$ extension more precisely[257, 258, 259, 260, 261]). Hence, one will expect the same condition for the preserved structure $\{I,\psi,\sigma\}\sim\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$, but $\sigma$ and $\sigma^{\prime}$ only satisfies the relation $h_{\sigma}+h_{\sigma^{\prime}}=0$. Hence, the flow $\{I,\psi,\sigma\}\rightarrow\{I^{\prime},\psi^{\prime},\sigma^{\prime}\}$ only preserves $Z_{2}$ structure when applying the anomaly classification to the non-group like symmetry whereas the flow $\{I,\psi,\sigma\}\rightarrow\{\overline{I}^{\prime},\overline{\psi}^{\prime},\overline{\sigma}^{\prime}\}$ is an anomaly preserving flow. More categorically, these two flows correspond to the choice of the common connected etale algebra $\{I,\psi\}=\{I^{\prime},\psi^{\prime}\}$ or $\{I,\psi,\sigma\}=\{\overline{I}^{\prime},\overline{\psi}^{\prime},\overline{\sigma}^{\prime}\}$. Consequently, if one believes that the massless flow $M(4,5)\rightarrow M(3,4)$ preserves the Ising fusion ring structure in the anomaly-free way, the application of the corresponding domain wall should change the chirality of objects. Phenomenologically, the domain wall changing the chiralities seems to correspond to the Alice ring in literature[151, 152].
@@ VIII Acknowledgement
The authors thank Po-Yao Chang, Naomichi Hatano, and Hosho Katsura for sharing their knowledge on non-Hermitian systems. YF thanks Yuma Furuta and Shinichiro Yahagi for related collaborations. YF also thanks Sylvain Ribault and Ingo Runkel for the discussions on fusion rules and Jurgen Fuch and Yuji Tachikawa for the helpful discussions on the distinction between fusion categories and fusion rings. We thank Hosho Katsura, Ingo Runkel and Gerard Watts for helpful comments and for checking some technical calculations in the manuscript. Also, we thank Tadashi Takayanagi for commenting on topological entanglement entropy and its importance, and Andrei Katsevich notifying us controvercies in the flow $M(2,7)\rightarrow M(2,5)$. YF thanks the support from NCTS. TK is supported by Grant-in-Aid for JSPS Fellows No. 23KJ1315.
@@ A.2 Phenomenology on entanglement cut: Naturality of complex spectrum
First thing to note is that the entanglement spectrum in a pseudo-Hermitian system can be complex[278, 279, 280, 281]. However, following the discussion in the previous sections, this phenomenon seems unusual because the complex energy value cannot appear in the previous discussion. Moreover, there exist established studies on the universal behaviors of real entanglement entropy in the nonunitary CFTs by analyzing the corresponding two dimensional statistical models and their quantum group structure[282, 283], combined with the twist field method[284, 285, 286]. We also note [287] as a reference summarizing a variant of related entropies, such as those in [288, 289] (for a recent analysis on symmetry resolved entanglement entropy[290, 291, 292, 293, 294, 295, 296, 297, 298] in nonunitary CFTs, see [299]).